# Present address: ]Lehrstuhl für Theoretische Chemie, Ruhr-Universität Bochum, 44780 Bochum, Germany

### v2

@@ I Introduction
The origin of life on Earth remains a profound scientific mystery. Numerous hypotheses have been put forward, but not a single one can explain all orgel1994origin ; mann2013origins ; sutherland2017opinion . As early as $\sim$150 years ago, Darwin proposed a “warm little pond” idea darwin1871jd , which was later developed by Oparin and Haldane into the influential “primordial soup” theory in the 1920s oparin1957origin , that is, the reactions of small molecules, such as CH4, NH3, H2O, and CO2, to form the first organic compounds. This leads to the abiotic origin of life on early Earth. In 1953, the famous Miller-Urey experiment was conducted to test this hypothesis at the presence of electric discharge miller1953production . Inspired by the “primordial soup” theory, various prebiotic environments have been proposed for the origin of life, ranging from the outer space ziurys2006chemistry ; sandford2020prebiotic ; krasnokutski2022pathway to deep-sea hydrothermal vents daniel2006origins ; martin2008hydrothermal .
@@
It is very challenging to study the origin of life in laboratory settings due to the extreme environments and subsequent analyses. Atomistic simulations based on first principles have been applied to study prebiotic chemistry, offering significant molecular-level insights perez2020prebiotic . Saitta et al. applied ab initio molecule dynamics (AIMD) and metadynamics simulations to study the Miller-Urey experiment, and found unexpected reaction intermediates including formic acid and formamide saitta2014miller . Goldman et al. applied AIMD simulations to investigate the formation of prebiotic compounds from impact-induced shock compression of cometary ices followed by expansion to ambient conditions, and found that the impact of cometary ices could produce amino acids independently of atmospheric conditions and Earth minerals goldman2010synthesis ; goldman2013prebiotic ; koziol2015prebiotic . Note that the shock compression duration was less than 10 ps, so their simulations did not reach thermal equilibrium, and it is unknown whether the biomolecules can remain stable without being quenched to ambient conditions. AIMD simulations coupled with enhanced sampling techniques have also been applied to study the role of mineral surfaces in prebiotic chemistry nair2006glycine ; pollet2006ab ; schreiner2011peptide . Despite the numerous previous studies, the exploration of prebiotic chemistry at both high hydrostatic pressures and high temperatures largely remains an uncharted territory.
@@ II Results and Discussion
Among these generated chemical products, formamide is a crucial precursor molecule that plays a vital role in the synthesis of biomolecules in prebiotic chemistry saladino2012formamide ; pietrucci2015formamide ; enchev2021chemical ; saladino2015meteorite ; rios2015impact ; ferus2015high . Thus, we first investigated the formation of formamide molecule. Fig. 3 (a) shows the different formation pathways at three P-T conditions, and detailed formation processes are displayed in Figure S4. At 10 GPa and 1000 K, the formation reaction begins with the dissociation of water molecules, generating OH- and H+ ions in the aqueous solution. The encounter between an H+ ion and a CO molecule leads to the formation of a C-H bond and an aldehyde group (-CH=O). Subsequently, the aldehyde group reacts with a NH3 molecule to generate a C-N bond, and then, one N-H bond breaks, ultimately producing formamide. At 10 GPa and 1400 K, the formation process differs from that at 10 GPa and 1000 K. One N-H bond in NH3 first breaks, leading to a dissociative amino group (-NH2). The amino group then reacts with a CO molecule, forming the C-N bond. Later, a water molecule dissociates, providing a proton that ultimately contributes to the formation of the formamide molecule. At 13 GPa and 1400 K, the C-N bond is initially formed through the reaction between NH3 and CO molecules. After a very short time, a dissociated proton from a water molecule is also bonded to the carbon atom, forming a C-H bond. Finally, by releasing a proton from the NH3 group, the formamide molecule is generated. Despite the different formation pathways at three P-T conditions, we observed two common features: first, the reactants always have CO, H2O, and NH3, and second, the reactions must involve the formation of C-N and C-H bonds and the breakage of N-H bonds.
@@
Previous studies showed that the cyclic forms of ribose are abundant in aqueous solutions at room temperature drew199813c . However, the cyclic forms of ribose were not detected in our simulations (Fig. 5 (a)). The limited size of our simulation boxes may restrict the configuration changes of ribose, thereby decreasing the likelihood of the C4’ or C5’ hydroxyl group attacking the C1’ aldehyde group. This limitation is unfavorable for the production of the furanose form (five-membered ring) or pyranose form (six-membered ring). To further study the cyclic formation mechanism, we carried a few additional simulations and finally chose the curved open-chain ribose molecules obtained from our previous simulations, as it appears that the C4’ or C5’ hydroxyl group has a greater chance of reacting with the C1’ aldehyde group under pressure. We built a new simulation box containing three of these ribose molecules together with 18 water molecules. After about 36 ps, we found the direct formation of a furanose form of ribose. Fig. 5 (c) shows the mechanism of ring formation. In this process, the C4’ hydroxyl group first loses a hydrogen atom, leaving a carbonyl group. The oxygen atom then attacks C1’ to form the ring, and finally, the carbonyl group on C1’ receives a dissolved proton in the solution to generate the five-membered ribose ring. Interestingly, no six-membered ring was found in our simulations at 10 GPa and 1400 K.
@@
Note that at ambient conditions, the naturally-occurring ribose exists as a mixture of cyclic forms in equilibrium with its open-chain form in aqueous solutions, and the pyranose form of ribose is the predominant conformation, accounting for approximately 70% dewick2013essentials ; dass2021equilibrium . However, despite the thermodynamic stability of the pyranose form of ribose over the furanose form at ambient conditions, the exclusive constituent of the carbohydrate backbone of RNA remains the furanose form of ribose. The reason behind it is still unknown. There are several hypotheses proposed to explain this preference, including silicate(borate)/ribose complexes kolb2004complexes ; lambert2004silicate and temperature gradients dass2021equilibrium . Our finding suggests that extreme P-T conditions may also generate the necessary component for assembling RNA molecules. Consequently, it highlights the potential role of upper mantle environment in facilitating the emergence and development of RNA-based life.
@@
We then shifted our attention to the chemical evolution of nucleobases, which are another essential building blocks of genetic information. There are five naturally-occurring nucleobases: adenine (A), cytosine (C), guanine (G), thymine (T), and uracil (U). Of these, uracil is unique to RNA, so our studies primarily focused on investigating the formation of uracil molecules under extreme conditions. We first studied the formation of the open-chain “C4N2” molecule which has the same CN-backbone as the open-chain form of uracil, as shown in Fig. 6 (a). Previous experimental studies on the origins of life have suggested that formamide is a crucial molecule for the prebiotic formation of nucleobases, in the presence of mineral or metal oxide catalysts saladino2012formamide , heatingenchev2021chemical , and high-energy free radicals ferus2015high ; ferus2014high , so here we also first chose formamide as a reactant. However, after about 144 ps, we did not observe the formation of the “C4N2” molecule at 10 GPa and 1400 K, suggesting that it may be difficult for small molecules to directly synthesize uracil in the AIMD simulations (Figure S9). Later, we chose the chemical reactants that have closer structural similarities to the “C4N2” molecule. The reactant molecules, including “C4N”, “C3N2”, “C2N2”, and “C3”, were generated from our first-stage AIMD simulations. A recent study suggested that uracil could be synthesized from the reactions between urea and “C3” molecules choe2022mechanism , so we specifically selected the C3H4O4 molecule (number 10 in Figure S5) and the urea molecule (CH4N2O, number 11 in Fig. 1) as reactants. The simulation results are presented in Figure S10. However, regardless of the inclusion of water molecules or the application of different pressure conditions, we did not observe any “C4N2” molecules with the CN-backbone of the open-chain form of uracil. Then we tried to find another pathways and conducted six more AIMD simulations to investigate the possibility of forming uracil through reactions involving above-mentioned molecules at various pressures and 1400 K (see Figures S11, S12 and S14).
@@
Later we chose allophanic acid (C2H4N2O3, number 13 in Fig. 1 ) and (Z)-ethene-1,2-diol (C2H4O2, number 7 in Figure S5) as reactants, and built a new simulation box containing 5 C2H4N2O3, 5 C2H4O2, and 10 H2, as shown in Fig.6 (c). At $\sim$ 10 GPa and 1400 K, a “C5N2” molecule, 3-(3-carboxyureido)-2,3-dihydroxypropanoic acid, was formed at around 28.8 ps, This molecule has the same CN-backbone as the open-chain form of thymine (T). As the simulation continued, a “C4N2” molecule, 2,3-dihydroxy-3-ureidopropanoic acid, labeled as “C4N2-2”, was formed, which has the same CN-backbone as the open-chain form of uracil. The formation process of these two kinds of molecules are shown in Figure S15. We further performed a enhanced sampling simulation to analyze the ring formation process at $\sim$ 10 GPa and 1400 K. The simulation box has one “C4N2-2” molecule and 32 water molecules, as shown in Fig. 6 (c). Finally, a uracil-like molecule, called uracil glycol, was synthesized, which has a uracil ring. The free energy curve in Fig. 6 (c) shows that the cyclic form is more stable than the open-chain form, indicating that the uracil-like molecule is possible to form in geological C-H-O-N fluids at extreme P-T conditions. The uracil-like molecule can then serve as a precursor for the synthesis of other three types of nucleobases, providing the necessary components for forming RNA molecules.
@@
Geochemical evidence of biotic activity suggests that the emergence of life can be traced back to as early as 3.8 billion years ago mojzsis1996evidence . The modern style of plate tectonics may have started about 3.2 billion years ago brenner2020paleomagnetic , but some form of plate tectonics may have already been active around 4.0 billion years ago. and these tectonic activities might carry water, carbon, and other volatile substances down into the mantle korenaga2013initiation ; HOLDER2024 ; blichert2008hafnium ; hopkins2008low ; huang2022barium . Because of reducing environments in early Earth’s mantle aulbach2016evidence , oxidized carbon and nitrogen can be reduced to CO and NH3/NH+${}_{4}^{+}$, respectively holm2009reduction (see Figure S16), and they may further synthesize to become key building blocks of life, which may reach the surface of early Earth. Our findings also have important implications for organic molecules in interstellar space. The reactant molecules H2O, H2, CO and NH3 largely exist in interstellar clouds and circumstellar shells, together with more than 200 complex organic species, such as aldehydes, alcohols, acids, amines and carboxamides guelin2022organic . It is unclear how those complex organic species are generated. Our current study suggests that the extreme pressure and temperature inside giant planets and their moons neri2020carbonaceous ; nettelmann2016uranus , like Jupiter and Saturn, may help to synthesize those complex organic molecules postberg2018macromolecular ; waite2009liquid ; tosi2024salts .
@@ III Conclusions
Extensive AIMD simulations have been performed to investigate the chemical reactions of small proto-organic molecules under the extreme P-T conditions. Three corresponding P-T systems were considered, i.e., 10 GPa - 1000 K, 10 GPa - 1400 K, and 13 GPa - 1400 K. The simulation results have revealed the potential formation of substantial quantities of organic matter without any catalysts, with the composition and abundance of CN-containing species being profoundly influenced by the specific P-T conditions. Notably, formamide, recognized as a crucial precursor in the production of biomolecules, is observed to readily form under extreme P-T conditions, and its formation pathways are compared across different P-T conditions. Furthermore, free energy calculations have confirmed that the 10 GPa - 1400 K condition is particularly favorable for the generation and preservation of CN-containing compounds.
Using the molecules formed, we have further explored the formation process of large biomolecules that are directly related to the building blocks of life. Remarkably, glycine, ribose, urea, and uracil-like molecules are successfully synthesized under extreme conditions in our AIMD simulations, thereby providing initial insights into the formation of organic compounds critical to the emergence of life. More importantly, different from that the six-membered ring of ribose is the majority in solutions at ambient conditions, it is found that formation of five-membered ring form of ribose is more favored at extreme P-T conditions, which could provide the necessary components for the generation of RNA molecules. Moreover, these compounds have the potential to be transported to Earth’s surface through geological activities, such as mantle convection, consequently contributing to the origins of life. The complex organic species found in interstellar space may be also related to the chemical processes inside giant planets and their moons at extreme P-T conditions. Our comprehensive studies present a novel perspective extending the cradle of life planetary interiors. The findings underscore the profound impact of extreme P-T conditions on the formation of organic matter and provide valuable insights into the mechanisms underlying the emergence of life within planetary bodies.
@@ III.1 Ab initio molecular dynamics simulations (AIMD)
We carried out AIMD simulations with the Born-Oppenheimer approximation using the Qbox code gygi2008architecture . We applied the PBE exchange-correlation functional in density functional theory perdew1996generalized . It is known that the PBE functional may be insufficient to describe aqueous solutions at ambient conditions, but our previous studies found that it works better at extreme P-T conditions than at ambient conditions pan2013dielectric ; pan2016fate ; stolte2019large ; stolte2021water , and many other studies also showed that this semi-local functional is reliable to study matter under extreme conditions spanu2011stability ; cheng2023thermodynamics ; zhang2023pbe . We used the ONCV pseudopotentials hamann2013optimized , and the Bussi-Donadio-Parrinello (BDP) thermostat, with a relaxation time of 24.2 fs, to control the temperature bussi2007canonical . To facilitate a larger time step of 0.24 fs, deuterium atoms were used in place of hydrogen atoms in our simulations. The simulation boxes have periodic boundary conditions. We first performed NPT simulations to achieve the desired pressures, configuring the plane-wave energy cutoff to 85 Ry and the energy cutoff for stress confinement potential to 70 Ry. We applied the simulation box sizes (Table S1) obtained from NPT simulations in the following NVT simulations for production runs, where the plane-wave energy cutoff was 65 Ry. During NVT simulations, we selected random time points to carry out independent simulations with energy cutoff of 85 Ry to monitor pressure (Figure S2). Our total AIMD simulation time exceeds 2.5 ns.
@@
To compute the free-energy landscape, we employed the adaptive biasing force method (ABF) darve2008adaptive as implemented in the software package SSAGES sidky2018ssages coupled with Qbox code. Unlike other advanced sampling methods that introduce biases into the energy, this method aims to flatten the generalized force. The running average of the force in the $k$th bin along the reaction coordinate $\xi$ is calculated as:
@@
where $\nu\left({{N}_{step}},k\right)$ is the number of samples collected in the $k$th bin after $N_{step}$ steps, ${{F}_{i}}\left(t_{i}^{k}\right)$ is the $i$th force sample at the time $t_{i}^{k}$, M is the mass matrix. The bias force -$\text{-}\left(\nabla\xi\right){{F}_{\xi}}\left({{N}_{step,}}k\right)$ is added to the calculation until the free energy landscape is flat. We finally integrated the total biased force to get the free energy. For example, the free energy $\Delta{{A}_{a\to b}}$ between two states $a$ and $b$ is calculated as:
@@
Supporting information
Details of the simulations, the percentage of C-N containing species over time, the hydrocarbon products, the free energy calculations, other strategies for the formation of ribose and uracil, and their ring structures, compositions of fluids at different oxygen fugacity.

### v3

@@ I Introduction
The origin of life on Earth remains a profound scientific mystery. Numerous hypotheses have been put forward, but not a single one can explain all orgel1994origin ; mann2013origins . As early as $\sim$150 years ago, Darwin proposed a “warm little pond” idea darwin1871jd , which was later developed by Oparin and Haldane into the influential “primordial soup” theory in the 1920s oparin1957origin , that is, the reactions of small molecules, such as CH4, NH3, H2O, and CO2, to form the first organic compounds. This leads to the abiotic origin of life on early Earth. In 1953, the famous Miller-Urey experiment was conducted to test this hypothesis at the presence of electric discharge miller1953production . Inspired by the “primordial soup” theory, various prebiotic environments have been proposed for the origin of life, ranging from the outer space ziurys2006chemistry ; sandford2020prebiotic ; krasnokutski2022pathway to deep-sea hydrothermal vents daniel2006origins ; martin2008hydrothermal .
@@
It is very challenging to study the origin of life in laboratory settings due to the extreme environments and subsequent analyses. Atomistic simulations based on first principles have been applied to study prebiotic chemistry, offering significant molecular-level insights perez2020prebiotic . Saitta et al. applied ab initio molecule dynamics (AIMD) and metadynamics simulations to study the Miller-Urey experiment, and found unexpected reaction intermediates including formic acid and formamide saitta2014miller . Goldman et al. applied AIMD simulations to investigate the formation of prebiotic compounds from impact-induced shock compression of cometary ices followed by expansion to ambient conditions, and found that the impact of cometary ices could produce amino acids independently of atmospheric conditions and Earth minerals goldman2010synthesis ; goldman2013prebiotic ; koziol2015prebiotic . Note that the shock compression simulations did not reach thermal equilibrium, and it is unknown whether the biomolecules can remain stable without being quenched to ambient conditions Ross2013comment ; Goldman2013reply . AIMD simulations coupled with enhanced sampling techniques have also been applied to study the role of mineral surfaces in prebiotic chemistry nair2006glycine ; pollet2006ab ; schreiner2011peptide . Despite the numerous previous studies, the exploration of prebiotic chemistry at both high hydrostatic pressures and high temperatures largely remains an uncharted territory.
@@ II Results and Discussion
Among these generated chemical products, formamide is a crucial precursor molecule that plays a vital role in the synthesis of biomolecules in prebiotic chemistry saladino2012formamide ; pietrucci2015formamide ; saladino2015meteorite ; rios2015impact ; ferus2015high ; enchev2021chemical . Thus, we first investigated the formation of formamide molecule. Fig. 3 (a) shows the different formation pathways at three P-T conditions, and detailed formation processes are displayed in Figure S4. At 10 GPa and 1000 K, the formation reaction begins with the dissociation of water molecules, generating OH- and H+ ions in the aqueous solution. The encounter between an H+ ion and a CO molecule leads to the formation of a C-H bond and an aldehyde group (-CH=O). Subsequently, the aldehyde group reacts with a NH3 molecule to generate a C-N bond, and then, one N-H bond breaks, ultimately producing formamide. At 10 GPa and 1400 K, the formation process differs from that at 10 GPa and 1000 K. One N-H bond in NH3 first breaks, leading to a dissociative amino group (-NH2). The amino group then reacts with a CO molecule, forming the C-N bond. Later, a water molecule dissociates, providing a proton that ultimately contributes to the formation of the formamide molecule. At 13 GPa and 1400 K, the C-N bond is initially formed through the reaction between NH3 and CO molecules. After a very short time, a dissociated proton from a water molecule is also bonded to the carbon atom, forming a C-H bond. Finally, by releasing a proton from the NH3 group, the formamide molecule is generated. Despite the different formation pathways at three P-T conditions, we observed two common features: first, the reactants always have CO, H2O, and NH3, and second, the reactions must involve the formation of C-N and C-H bonds and the breakage of N-H bonds.
@@
Previous studies showed that the cyclic forms of ribose are abundant in aqueous solutions at room temperature drew199813c . However, the cyclic forms of ribose, either furanose form (five-membered ring) or pyranose form (six-membered ring), were not detected in our simulations (Fig. 5 (a)). The configuration change of ribose may require very long simulation time. To further study the cyclic formation mechanism, we tried a few additional simulations with new initial states, and finally chose the curved open-chain ribose molecules obtained from our previous simulations, as it appears that the C4’ or C5’ hydroxyl group has a greater chance of reacting with the C1’ aldehyde group under pressure. We built a new simulation box containing three of these ribose molecules together with 18 water molecules. After about 36 ps, we observed the direct formation of a furanose form of ribose. Fig. 5 (c) shows the mechanism of ring formation. In this process, the C4’ hydroxyl group first loses a hydrogen atom, leaving a carbonyl group. The oxygen atom then attacks C1’ to form the ring, and finally, the carbonyl group on C1’ receives a dissolved proton in the solution to generate the five-membered ribose ring. Interestingly, no six-membered ring was found in our simulations at 10 GPa and 1400 K.
@@
Note that at ambient conditions, the naturally-occurring ribose exists as a mixture of cyclic forms in equilibrium with its open-chain form in aqueous solutions, and the pyranose form of ribose is the predominant conformation, accounting for approximately 70% dewick2013essentials ; dass2021equilibrium . However, despite the thermodynamic stability of the pyranose form of ribose over the furanose form at ambient conditions, the exclusive constituent of the carbohydrate backbone of RNA remains the furanose form of ribose. The reason behind it is still unknown. There are several hypotheses proposed to explain this preference, including silicate(borate)/ribose complexes kolb2004complexes ; lambert2004silicate and temperature gradients dass2021equilibrium . Our finding suggests that extreme P-T conditions may also generate the necessary component for assembling RNA molecules. Consequently, it highlights the potential role of deep Earth environment in facilitating the emergence and development of RNA-based life.
@@
We then shifted our attention to the chemical evolution of nucleobases, which are another essential building blocks of genetic information. There are five naturally-occurring nucleobases: adenine (A), cytosine (C), guanine (G), thymine (T), and uracil (U). Of these, uracil is unique to RNA, so our studies primarily focused on investigating the formation of uracil molecules under extreme conditions. We first studied the formation of the open-chain “C4N2” molecule which has the same CN-backbone as the open-chain form of uracil, as shown in Fig. 6 (a). Previous experimental studies on the origin of life have suggested that formamide is a crucial molecule for the prebiotic formation of nucleobases, in the presence of mineral or metal oxide catalysts saladino2012formamide , heatingenchev2021chemical , and high-energy free radicals ferus2015high ; ferus2014high , so here we also first chose formamide as a reactant. However, after about 144 ps, we did not observe the formation of the “C4N2” molecule at 10 GPa and 1400 K, suggesting that it may be difficult for small molecules to directly synthesize uracil in the AIMD simulations (Figure S9). Later, we chose the chemical reactants that have closer structural similarities to the “C4N2” molecule. The reactant molecules, including “C4N”, “C3N2”, “C2N2”, and “C3”, were generated from our first-stage AIMD simulations. A recent study suggested that uracil could be synthesized from the reactions between urea and “C3” molecules choe2022mechanism , so we specifically selected the C3H4O4 molecule (number 10 in Figure S5) and the urea molecule (CH4N2O, number 11 in Fig. 1) as reactants. The simulation results are presented in Figure S10. However, regardless of the inclusion of water molecules or the application of different pressure conditions, we did not observe any “C4N2” molecules with the CN-backbone of the open-chain form of uracil. Then we tried to find another pathways and conducted six more AIMD simulations to investigate the possibility of forming uracil through reactions involving above-mentioned molecules at various pressures and 1400 K (see Figures S11, S12 and S14).
@@
Later we chose allophanic acid (C2H4N2O3, number 13 in Fig. 1 ) and (Z)-ethene-1,2-diol (C2H4O2, number 7 in Figure S5) as reactants, and built a new simulation box containing 5 C2H4N2O3, 5 C2H4O2, and 10 H2, as shown in Fig.6 (c). At $\sim$ 10 GPa and 1400 K, a “C5N2” molecule, 3-(3-carboxyureido)-2,3-dihydroxypropanoic acid, was formed at around 28.8 ps, This molecule has the same CN-backbone as the open-chain form of thymine (T). As the simulation continued, a “C4N2” molecule, 2,3-dihydroxy-3-ureidopropanoic acid, labeled as “C4N2-2”, was formed, which has the same CN-backbone as the open-chain form of uracil. The formation process of these two kinds of molecules are shown in Figure S15. We further performed an enhanced sampling simulation to analyze the ring formation process at $\sim$ 10 GPa and 1400 K. The simulation box has one “C4N2-2” molecule and 32 water molecules, as shown in Fig. 6 (c). Finally, a uracil-like molecule, called uracil glycol, was synthesized, which has a uracil ring. The free energy curve in Fig. 6 (c) shows that the cyclic form is more stable than the open-chain form, indicating that the uracil-like molecule is possible to form in geological C-H-O-N fluids at extreme P-T conditions. The uracil-like molecule can then serve as a precursor for the synthesis of other three types of nucleobases, providing the necessary components for forming RNA molecules.
@@
Geochemical evidence of biotic activity suggests that the emergence of life can be traced back to as early as 3.8 billion years ago mojzsis1996evidence . The modern style of plate tectonics may have started about 3.2 billion years ago brenner2020paleomagnetic , but some form of plate tectonics may have already been active around 4.0 billion years ago, and these tectonic activities might carry water, carbon, and other volatile substances down into the mantle korenaga2013initiation ; HOLDER2024 ; blichert2008hafnium ; hopkins2008low ; huang2022barium . Because of reducing environments in early Earth’s mantle aulbach2016evidence , oxidized carbon and nitrogen can be reduced to CO and NH3/NH+${}_{4}^{+}$, respectively (see Figure S16)holm2009reduction , and they may further synthesize to become key building blocks of life, which may reach the surface of early Earth. Our findings also have important implications for organic molecules in interstellar space. The reactant molecules H2O, H2, CO and NH3 largely exist in interstellar clouds and circumstellar shells, together with more than 200 complex organic species, such as aldehydes, alcohols, acids, amines and carboxamides guelin2022organic . It is unclear how those complex organic species are generated. Our study suggests that the extreme P and T inside giant planets and their moons neri2020carbonaceous ; nettelmann2016uranus , like Jupiter and Saturn, may help to synthesize those complex organic molecules postberg2018macromolecular . Earth’s or other giant planets’ interiors also provide a protective environment that filters out harmful radiation and buffers against drastic physico-chemical variations.
@@ III Conclusions
Extensive AIMD simulations have been performed to investigate the chemical reactions of small proto-organic molecules under the extreme P-T conditions. Three corresponding P-T systems were considered, i.e., 10 GPa - 1000 K, 10 GPa - 1400 K, and 13 GPa - 1400 K. The simulation results revealed the potential formation of substantial quantities of organic matter without any catalysts, with the composition and abundance of CN-containing species being profoundly influenced by the specific P-T conditions. Notably, formamide, recognized as a crucial precursor in the production of biomolecules, is observed to readily form under extreme P-T conditions, and its formation pathways are compared across different P-T conditions. Furthermore, free energy calculations confirmed that the 10 GPa - 1400 K condition is particularly favorable for the generation and preservation of CN-containing compounds.
Using the molecules formed, we have further explored the formation process of large biomolecules that are directly related to the building blocks of life. Remarkably, glycine, ribose, urea, and uracil-like molecules are successfully synthesized under extreme conditions in our AIMD simulations, thereby providing initial insights into the formation of organic compounds critical to the emergence of life. Particularly, different from that the six-membered ring form of ribose is the majority in solutions at ambient conditions, it was found that the formation of the five-membered ring form of ribose is more favored at extreme P-T conditions, which provides the necessary component for the generation of RNA molecules. Moreover, these compounds have the potential to be transported to Earth’s surface through geological activities, such as mantle convection, consequently contributing to the origin of life. The complex organic species found in interstellar space may be also related to the chemical processes inside giant planets and their moons at extreme P-T conditions. Our comprehensive studies present a novel perspective extending the cradle of life in planetary interiors. The findings underscore the profound impact of extreme P-T conditions on the formation of organic matter and provide valuable insights into the mechanisms underlying the emergence of life within planetary bodies.
@@ III.1 Ab initio molecular dynamics simulations (AIMD)
We carried out AIMD simulations with the Born-Oppenheimer approximation using the Qbox code gygi2008architecture . We applied the PBE exchange-correlation functional in density functional theory perdew1996generalized . It is known that the PBE functional may be insufficient to describe aqueous solutions at ambient conditions, but our previous studies found that it works better at extreme P-T conditions than at ambient conditions pan2013dielectric ; pan2016fate ; stolte2019large ; stolte2021water , and many other studies also showed that this semi-local functional is reliable to study matter under extreme conditions spanu2011stability ; cheng2023thermodynamics ; zhang2023pbe . We used the ONCV pseudopotentials hamann2013optimized , and the Bussi-Donadio-Parrinello (BDP) thermostat, with a relaxation time of 24.2 fs, to control the temperature bussi2007canonical . To facilitate a larger time step of 0.24 fs, deuterium atoms were used in place of hydrogen atoms in our simulations. The simulation boxes have periodic boundary conditions. We first performed NPT simulations to achieve the desired pressures, configuring the plane-wave kinetic energy cutoff to 85 Ry and the cutoff for stress confinement potential to 70 Ry focher1994structural . We applied the simulation box sizes (Table S1) obtained from NPT simulations in the following NVT simulations for production runs, where the plane-wave kinetic energy cutoff was 65 Ry. During NVT simulations, we selected random time points to carry out independent simulations with the cutoff of 85 Ry to monitor pressure (Figure S2). Our total AIMD simulation time exceeds 2.5 ns.
@@
To compute the free-energy landscape, we employed the adaptive biasing force method (ABF) darve2008adaptive as implemented in the software package SSAGES sidky2018ssages coupled with Qbox code. Unlike other advanced sampling methods that introduce bias potentials, this method aims to flatten the generalized force. The running average of the force in the $k$th bin along the reaction coordinate $\xi$ is calculated as:
@@
where $\nu\left({{N}_{step}},k\right)$ is the number of samples collected in the $k$th bin after $N_{step}$ steps, ${{F}_{i}}\left(t_{i}^{k}\right)$ is the $i$th force sample at the time $t_{i}^{k}$, and M is the mass matrix. The bias force -$\text{-}\left(\nabla\xi\right){{F}_{\xi}}\left({{N}_{step,}}k\right)$ is added to the calculation until the free energy landscape is flat. We finally integrated the total biased force to get the free energy. For example, the free energy $\Delta{{A}_{a\to b}}$ between two states $a$ and $b$ is calculated as: